OpenStudy (anonymous):

Which of the following is the solution to the equation 8^(y – 2) = 2^y ? y = 1 y = 3 y = –3 ---> My Answer y = –1

OpenStudy (anonymous):

@bradely

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

First step...rewrite the left side like this..since 8 is just 2^3 \[2^{3(y-2)}=2^{y}\]

OpenStudy (anonymous):

Now you can just set 3(y-2) = y since you have the same base (2) on both sides

OpenStudy (anonymous):

You shouldn't get -3

OpenStudy (anonymous):

i dont understand this

OpenStudy (anonymous):

You dont know what to do next? Or you don't understand one of my steps?

OpenStudy (anonymous):

your steps actually helped me but i dont understand what to do next i would ask my teacher but she gets mad when i ask her to explain things to me

OpenStudy (anonymous):

Like I said a few lines up, just solve 3(y-2) = y